### v2

@@ D.1 Motivation
For what purpose was the dataset created? The dataset was created to address the current limitations of a lack of high-fidelity training data for the development and testing of machine learning methods for Computational Fluid Dynamics. In addition, it was created to be used as a dataset for the 4th Automotive CFD Prediction Workshop 666https://autocfd.org
@@ D.2 Distribution
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? The dataset will be free to download from Amazon S3 (without the need for an AWS account).
When will the dataset be distributed? The dataset is already available to download via Amazon S3.
@@
D.3 Maintenance
Who will be supporting/hosting/maintaining the dataset? AWS are hosting the dataset on Amazon S3
How can the owner/curator/manager of the dataset be contacted (e.g., email address)? The owner/curator/manager of the dataset can be contacted at neashton@amazon.com (these are also provided in the dataset README and paper).
Is there an erratum? No, but if we find errors we will provide updates to the dataset and note any changes in the dataset README.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? Yes the dataset will be updated to address errors or provided extra functionality. The README of the dataset will be updated to reflect this.
@@ D.3 Maintenance
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? We will consider this on a case by case basis and they can contact neashton@amazon.com to discuss this further.
D.4 Composition
@@
Are there recommended data splits (e.g., training, development/validation, testing)? We don’t provide explicit recommendations.
@@ D.5 Collection Process
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? The main paper discusses the HPC hardware used
@@
D.6 Preprocessing/cleaning/labeling

### v3

@@ E.1 Motivation
For what purpose was the dataset created? The dataset was created to address the current limitations of a lack of high-fidelity training data for the development and testing of machine learning methods for Computational Fluid Dynamics. In addition, it was created to be used as a dataset for the 4th Automotive CFD Prediction Workshop 888https://autocfd.org
@@ E.2 Distribution
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? The dataset is free to download from Amazon S3 (without the need for an AWS account) and is fully described on the dataset website 999https://caemldatasets.org. Additional download sites and options are in the process of being created and will be shared on the website once ready.
When will the dataset be distributed? The dataset is already available to download via Amazon S3. Additional download sites and options are in the process of being created and will be shared on the website once ready.
@@
E.3 Maintenance
Who will be supporting/hosting/maintaining the dataset? The dataset is being managed by the collection of authors and a public website 101010https://caemldatasets.org will provide on-going updates on hosting/maintenance.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)? The owner/curator/manager of the dataset can be contacted at contact@caemldatasets.org (these are also provided in the dataset README and paper).
Is there an erratum? No, but if we find errors we will provide updates to the dataset and note any changes in the dataset README and website.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? Yes the dataset will be updated to address errors or provided extra functionality. The README of the dataset and the dataset website will be updated to reflect this.
@@ E.3 Maintenance
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? We will consider this on a case by case basis and they can contact contact@caemldatasets.org to discuss this further.
E.4 Composition
@@
Are there recommended data splits (e.g., training, development/validation, testing)? 60/20/20 is the recommended split based upon initial testing.
@@ E.5 Collection Process
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensors, manual human curation, software programs, software APIs The main paper discusses the HPC hardware used
@@
E.6 Preprocessing/cleaning/labeling